OpenStudy (anonymous):

I wil fan and medal! Solve three-fourths z minus one-sixth equals negative one-half for z. Choose one answer. A. z equals negative four B. z equals negative eight-ninths C. z equals negative four-ninths D. z equals negative one-fourth

OpenStudy (anonymous):

@technodos

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

(3/4)z - (1/6) = -(1/2) is this the problem?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

ok, what do you get when you add 1/6 to each side?

OpenStudy (anonymous):

ok what is -1/2 +1/6

Vocaloid (vocaloid):

hint: -1/2 = -3/6

OpenStudy (anonymous):

ok

Vocaloid (vocaloid):

so, what is -3/6+1/6?

OpenStudy (anonymous):

-2/6

Vocaloid (vocaloid):

yes! so, now, what is -1/2 + 1/6?

OpenStudy (anonymous):

-2/6

OpenStudy (anonymous):

what's 3/4 z=-2/6

OpenStudy (anonymous):

hello? i need..help!

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@GrimVix

OpenStudy (anonymous):

@h

OpenStudy (anonymous):

@jordanjamesbay

OpenStudy (anonymous):

@technodos

OpenStudy (anonymous):

@technodos

OpenStudy (anonymous):

HELP!!!!

OpenStudy (anonymous):

what is -2/6 divided by 3/4

Vocaloid (vocaloid):

(-2/6)/(3/4) = (-2/6)*(4/3) to divide fractions, multiply the first fraction by the reciprocal of the second fraction

OpenStudy (anonymous):

Choose one answer. A. z equals negative four B. z equals negative eight-ninths C. z equals negative four-ninths D. z equals negative one-fourth

OpenStudy (anonymous):

it doesn't match my answers

OpenStudy (texaschic101):

There are a couple of ways to do this... 3/4z - 1/6 = - 1/2......to get rid of fractions, multiply everything by the common denominator of 12 (12)3/4z - (12)1/6 = (12) - 1/2 9z - 2 = - 6 --- add 2 to both sides 9z = -6 + 2 9z = -4 z = -4/9 OR THIS WAY 3/4z - 1/6 = -1/2 --- add 1/6 to both sides 3/4z = -1/2 + 1/6 3/4z = - 3/6 + 1/6 3/4z = -2/6 -- divide both sides by 3/4 z = (-2/6) / (3/4) z = -2/6 * 4/3 z = - 8/18 which reduces to - 4/9

OpenStudy (texaschic101):

understand ?